Suppose that the daily income of Mundungundunguzika rises from K50 to K100. As a result of...

  1. Suppose that the daily income of Mundungundunguzika rises from K50 to K100. As a result of the increase, his purchases of bread increases from 2 loaves per day to 3 loaves per day. What is his income elasticity of demand for bread? [3 Marks]
  2. Suppose now that Mundungundunguzika faces the following demand function

ANvz8ZJySBLQAAAAAElFTkSuQmCC

Where Y is his income, Q is the quantity of Bread purchases and P is the price of bread per loaf.

  1. When Price is K150 and income is K15,000, determine the following:
  1. Price elasticity of demand [2 Marks]
  2. Income elasticity of demand [2 Marks]
  1. Determine what effect a rise in price would have on total revenue. [1 Mark]
  2. Assess how sales of bread would change during a period of rising incomes               [2 Marks]
Answer #1

Now, the income increasis from a encome increasybrom k50 to kloo. pm clase of bread is ceases from 2 loaus Pu and day to 3 loi suppose how that mundug unduguzika faces the following demand function ro=(5 4000 + 30y-29 ) 4000 H307-2Q 2) P = 4 od +30y&= -5.5A 1150 226175 = -0.0036 the absolate value of elasticity -0.004 lel=0.004 ay Now Q = -5.50 +2000 tisy Now O Q = 15 Now